OpenStudy (anonymous):

A thief 40 meters away runs towards you at 8m/s. What is the smallest acceleration so that v=at keeps you in front? The answer is 8/5 m/s^2. However, I have tried 1 m/s^2 (which is smaller than the answer) and it still can keep the person in front... Many thanks!

OpenStudy (kropot72):

The thief's target must reach a speed of 8 m/s in t seconds. Assuming constant acceleration, the average speed of the target during acceleration is 8/2 = 4 m/s. The distance traveled by the target during the acceleration period is 4t. The distance traveled by the thief while his target is accelerating is (40 + 4t) m. The time taken by the thief to just reach his target is: \[\frac{40+4t}{8}=5+0.5t\ seconds\] But this time is equal to the time for the target to accelerate to 8 m/s. Therefore: \[5+0.5t=t\] Solving gives the value for t as 10 seconds. Substituting in v = at we get: 8 = 10a Therefore the minimum acceleration is 0.8 m/s^2